In a class of 15 people exactly 3 got A's. If two people are randomly chosen from this class, what is the probability that at least one of these two people got and A?
Answer by stanbon(75887)

In a class of 15 people exactly 3 got A's. If two people are randomly chosen from this class, what is the probability that at least one of these two people got and A?
-----
P(at least 1 got A) = 1 - P(none got A)
-----
P(none got A) = 12C2/15C2 = 0.6286
-----------------
P(at least 1 got A) = 1 - 0.6286 = 0.3714
===============================================
Cheers,
Stan H.
===============